# mystyle backgroundcolor=white!10

### v1

@@ VI Conclusion
In this paper, we introduced novel sufficient conditions for the finite-time safety and reach-avoid verification of stochastic discrete-time dynamical systems. These conditions effectively determine the lower and upper bounds of the probability that, within a predefined finite-time horizon, a system starting from an initial state in a safe set will either exit the safe set (safety verification) or reach a target set while remaining within the safe set until the first encounter with the target (reach-avoid verification). They complement existing criteria or bridge existing gaps in the literature. Finally, we demonstrated their performance in finite-time safety and reach-avoid verification on two numerical examples, utilizing semi-definite programming tools.

### v2

@@ 6 Conclusion
In this paper, we introduced novel sufficient conditions for the finite-time safety and reach-avoid verification of stochastic discrete-time dynamical systems. These conditions provide the lower and upper bounds of the probability that, within a predefined finite-time horizon, a system starting from an initial state in a safe set will either exit the safe set (safety verification) or reach a target set while remaining within the safe set until the first encounter with the target (reach-avoid verification). They complement existing criteria or bridge existing gaps in the literature. Finally, we demonstrated their performance in finite-time safety and reach-avoid verification on two numerical examples, utilizing semi-definite programming tools.
In the future, I would like to rigorously assess the conservativeness of the derived bounds through a necessity analysis of the proposed barrier-like conditions, extending the infinite-time framework developed in [33]. Furthermore, I would develop analysis methods for probabilistic programs by integrating these barrier-like conditions. For example, I intend to investigate termination analysis of probabilistic programs within bounded time horizons and compare the results with state-of-the-art approaches, such as those presented in [8].
Acknowledgements
This work is funded by the CAS Pioneer Hundred Talents Program, Basic Research Program of Institute of Software, CAS (Grant No. ISCAS-JCMS-202302), and NRF RSS Scheme NRF-RSS2022-009.